\begin{thebibliography}{99}
\bibitem{Pelaez:2015qba}
J.~R.~Pelaez,
%``From controversy to precision on the sigma meson: 
%a review on the status of the non-ordinary $f_0(500)$ resonance,''
Phys. Rept. \textbf{658}, 1 (2016). %[arXiv:1510.00653 [hep-ph]].

\bibitem{pdg}
S. Navas \textit{et al.} (Particle Data Group),
Phys. Rev. D \textbf{110}, 030001 (2024).


\bibitem{Pelaez:2020uiw}
J.~R.~Pel\'aez and A.~Rodas,
%``Determination of the lightest strange resonance $K_0^*(700)$ or $\kappa$, from a dispersive data analysis,''
Phys. Rev. Lett. \textbf{124}, 172001 (2020). %[arXiv:2001.08153 [hep-ph]].

\bibitem{Pelaez:2020gnd}
J.~R.~Pel\'aez and A.~Rodas, Phys. Rept. \textbf{969}, 1 (2022). %[arXiv:2010.11222 [hep-ph]].

% DtoKpipi: ===========
\bibitem{E791:2002xlc}
E.~M.~Aitala \textit{et al.} [E791],
%``Dalitz Plot Analysis of the Decay $D^+ \to K^- \pi^+ \pi^+$ and the Study of the $K\pi$ Scalar Amplitudes,''
Phys. Rev. Lett. \textbf{89}, 121801 (2002). %[arXiv:hep-ex/0204018 [hep-ex]].

\bibitem{FOCUS:2007mcb}
J.~M.~Link \textit{et al.} [FOCUS],
%``Dalitz plot analysis of the $D^{+} \to K^{-} \pi^{+} \pi^{+}$ decay in the FOCUS experiment,''
Phys. Lett. B \textbf{653}, 1 (2007). %[arXiv:0705.2248 [hep-ex]].

\bibitem{CLEO:2008jus}
G.~Bonvicini \textit{et al.} [CLEO], %``Dalitz plot analysis of the D+ ---\ensuremath{>} K- pi+ pi+ decay,''
Phys. Rev. D \textbf{78}, 052001 (2008). %[arXiv:0802.4214 [hep-ex]].

\bibitem{BESIII:2014oag}
M.~Ablikim \textit{et al.} [BESIII], %``Amplitude Analysis of the $D^+ \to K^0_S \pi^+ \pi^0$ Dalitz Plot,''
Phys. Rev. D \textbf{89}, 052001 (2014). %[arXiv:1401.3083 [hep-ex]].

% partial-wave: ========

\bibitem{FOCUS:2005iqy}
J.~M.~Link \textit{et al.} [FOCUS],
%``Hadronic mass spectrum analysis of D+ ---\ensuremath{>} K- pi+ mu+ nu decay 
%and measurement of the K*(892)0 mass and width,''
Phys. Lett. B \textbf{621}, 72 (2005). %[arXiv:hep-ex/0503043 [hep-ex]].

\bibitem{CLEO:2010enr}
R.~A.~Briere \textit{et al.} [CLEO],
%``Analysis of $D^+ \to K^- \pi^+ e^+ \nu_e$ and $D^+ \to K^- \pi^+ \mu^+ \nu_\mu$ Semileptonic Decays,''
Phys. Rev. D \textbf{81}, 112001 (2010). %[arXiv:1004.1954 [hep-ex]].

\bibitem{BaBar:2010vmf}
P.~del Amo Sanchez \textit{et al.} [BaBar],
%``Analysis of the $D^+ \to K^- \pi^+ e^+ \nu_e$ decay channel,''
Phys. Rev. D \textbf{83}, 072001 (2011). %[arXiv:1012.1810 [hep-ex]].

\bibitem{BESIII:2015hty}
M.~Ablikim \textit{et al.} [BESIII],
%``Study of $D^{+} \to K^{-} \pi^+ e^+ \nu_e$,''
Phys. Rev. D \textbf{94}, 032001 (2016). %[arXiv:1512.08627 [hep-ex]].

\bibitem{BESIII:2018jjm}
M.~Ablikim \textit{et al.} [BESIII],
%``Study of the decay $D^0\rightarrow \bar{K}^0\pi^-e^+\nu_e$,''
Phys. Rev. D \textbf{99}, 011103 (2019). %[arXiv:1811.11349 [hep-ex]].

\bibitem{BESIII:2024awg}
M.~Ablikim \textit{et al.} [BESIII],
%``Analysis of the dynamics of the decay $ {D}^{+}\to {K}_S^0{\pi}^0{e}^{+}{\nu}_e $,''
JHEP \textbf{10}, 199 (2024). %[arXiv:2408.04422 [hep-ex]].

\bibitem{BESIII:2024xjf}
M.~Ablikim \textit{et al.} [BESIII],
%``Study of the semileptonic decay $ {D}^0\to {\overline{K}}^0{\pi}^{-}{e}^{+}{\nu}_e $,''
JHEP \textbf{03}, 197 (2025). %[arXiv:2412.10803 [hep-ex]].

\bibitem{BESIII:2025fso}
M.~Ablikim \textit{et al.} [BESIII],
%``Observation of $D^+\to K^0_S\pi^0\mu^+\nu_\mu$, 
%Test of Lepton Flavor Universality and First Angular Analysis of $D^+\to \bar{K}^\ast(892)^0\ell^+\nu_\ell$,''
arXiv:2506.05761 [hep-ex].


% q^2\bar q^2: ===================================
\bibitem{Jaffe:1976ig}
R.~L.~Jaffe, %``Multi-Quark Hadrons. 1. The Phenomenology of (2 Quark 2 anti-Quark) Mesons,''
Phys. Rev. D \textbf{15}, 267 (1977).

\bibitem{Jaffe:1976ih}
R.~L.~Jaffe, %``Multi-Quark Hadrons. 2. Methods,''
Phys. Rev. D \textbf{15}, 281 (1977).

\bibitem{Close:2002zu}
F.~E.~Close and N.~A.~Tornqvist,
%``Scalar mesons above and below 1-GeV,''
J. Phys. G \textbf{28}, R249 (2002). %[arXiv:hep-ph/0204205 [hep-ph]].

\bibitem{Pelaez:2003dy}
J.~R.~Pelaez,
%``On the Nature of light scalar mesons from their large N(c) behavior,''
Phys. Rev. Lett. \textbf{92}, 102001 (2004). %[arXiv:hep-ph/0309292 [hep-ph]].


\bibitem{Maiani:2004uc}
L.~Maiani, F.~Piccinini, A.~D.~Polosa and V.~Riquer,
%``A New look at scalar mesons,''
Phys. Rev. Lett. \textbf{93}, 212002 (2004). %[arXiv:hep-ph/0407017 [hep-ph]].

\bibitem{Amsler:2004ps}
C.~Amsler and N.~A.~Tornqvist,
%``Mesons beyond the naive quark model,''
Phys. Rept. \textbf{389}, 61 (2004).

\bibitem{Jaffe:2004ph}
R.~L.~Jaffe, %``Exotica,''
Phys. Rept. \textbf{409}, 1 (2005). %[arXiv:hep-ph/0409065 [hep-ph]].

\bibitem{tHooft:2008rus}
G.~'t Hooft, G.~Isidori, L.~Maiani, A.~D.~Polosa and V.~Riquer,
%``A Theory of Scalar Mesons,''
Phys. Lett. B \textbf{662}, 424 (2008). %[arXiv:0801.2288 [hep-ph]].

\bibitem{Fariborz:2009cq}
A.~H.~Fariborz, R.~Jora and J.~Schechter, %``Global aspects of the scalar meson puzzle,''
Phys. Rev. D \textbf{79}, 074014 (2009). %[arXiv:0902.2825 [hep-ph]].

\bibitem{Weinberg:2013cfa}
S.~Weinberg, %``Tetraquark Mesons in Large $N$ Quantum Chromodynamics,''
Phys. Rev. Lett. \textbf{110}, 261601 (2013). %[arXiv:1303.0342 [hep-ph]].

\bibitem{Agaev:2018fvz} 
S.~S.~Agaev, K.~Azizi and H.~Sundu,
%``The nonet of the light scalar tetraquarks: The mesons $a_0(980)$ and $K_{0}^{\ast }(800)$,''
Phys.\ Lett.\ B {\bf 789}, 405 (2019). %[arXiv:1804.02519 [hep-ph]].

% Formalism: =========
\bibitem{Buchalla:1995vs}
G.~Buchalla, A.~J.~Buras and M.~E.~Lautenbacher,
%``Weak decays beyond leading logarithms,''
Rev. Mod. Phys. {\bf 68}, 1125 (1996). %[arXiv:hep-ph/9512380 [hep-ph]].

\bibitem{Buras:1998raa} %``Weak Hamiltonian, CP violation and rare decays,''
A.~J.~Buras, hep-ph/9806471.

\bibitem{Hsiao:2023qtk}
Y.~K.~Hsiao, S.~Q.~Yang, W.~J.~Wei and B.~C.~Ke,
%``Testing the light scalar meson as a non-$ q\overline{q} $ state in semileptonic D decays,''
JHEP \textbf{12}, 226 (2025). %[arXiv:2306.06091 [hep-ph]].

\bibitem{Lee:1992ih}
C.~L.~Y.~Lee, M.~Lu and M.~B.~Wise, %``B(l4) and D(l4) decay,''
Phys. Rev. D \textbf{46}, 5040 (1992).

\bibitem{Tsai:2021ota}
S.~Y.~Tsai and Y.~K.~Hsiao,
%``Semileptonic $B^{-}\to\pi^{+}\pi^{-}\ell^{-}\bar\nu_\ell$ decay with $\pi\pi$ invariant mass spectrum,''
arXiv:2107.03634 [hep-ph].


\bibitem{Melikhov:2000yu}
D.~Melikhov and B.~Stech, %``Weak form-factors for heavy meson decays: An Update,''
Phys. Rev. D \textbf{62}, 014006 (2000). %[arXiv:hep-ph/0001113 [hep-ph]].
% ==================
\bibitem{Cheng:2022vbw}
H.~Y.~Cheng, C.~W.~Chiang and Z.~Q.~Zhang,
%``Hadronic three-body D decays mediated by scalar resonances,''
Phys. Rev. D \textbf{105}, 033006 (2022). %[arXiv:2201.00460 [hep-ph]].

\bibitem{Cheng:2020ipp}
H.~Y.~Cheng and C.~K.~Chua,
%``Branching fractions and $CP$ violation in $B^-\to K^+K^-\pi^-$ and $B^-\to \pi^+\pi^-\pi^-$ decays,''
Phys.\ Rev.\ D {\bf 102}, 053006 (2020). %[arXiv:2007.02558 [hep-ph]].

\bibitem{Boito:2009qd}
D.~R.~Boito and R.~Escribano,
%``Kpi form-factors and final state interactions in D+ ---{\ensuremath{>}} K- pi+ pi+ decays,''
Phys. Rev. D \textbf{80}, 054007 (2009). %[arXiv:0907.0189 [hep-ph]].

\bibitem{Escribano:2023zjx}
R.~Escribano, P.~Masjuan and P.~Sanchez-Puertas,
%``Combined study of hadronic D+{\textrightarrow}K-{\ensuremath{\pi}}+{\ensuremath{\pi}}+ and Ds+{\textrightarrow}K+K+{\ensuremath{\pi}}- decays by means of the analysis of semileptonic D+{\textrightarrow}K-{\ensuremath{\pi}}+{\ensuremath{\ell}}+{\ensuremath{\nu}}{\ensuremath{\ell}} decays,''
Phys. Rev. D \textbf{108}, 094006 (2023). %[arXiv:2302.03312 [hep-ph]].

\bibitem{Ali:1998eb}
A.~Ali, G.~Kramer and C.~D.~Lu,
%``Experimental tests of factorization in charmless nonleptonic two-body B decays,''
Phys. Rev. D \textbf{58}, 094009 (1998). %[arXiv:hep-ph/9804363 [hep-ph]].

\bibitem{Hsiao:2014mua}
Y.~K.~Hsiao and C.~Q.~Geng,
%``Direct CP violation in $\Lambda_b$ decays,''
Phys. Rev. D \textbf{91}, 116007 (2015). %[arXiv:1412.1899 [hep-ph]].

\bibitem{Wang:2024mjw}
Y.~L.~Wang, Y.~K.~Hsiao, K.~L.~Wang and C.~C.~Lih,
%``Charmful two-body {\ensuremath{\Omega}}b decays in the light-front quark model,''
Phys. Rev. D \textbf{111}, 096013 (2025). %[arXiv:2412.11584 [hep-ph]].

% ================
\bibitem{Branz:2007xp}
T.~Branz, T.~Gutsche and V.~E.~Lyubovitskij,
%``f0(980) meson as a K anti-K molecule in a phenomenological Lagrangian approach,'' (f0-->KKbar)
Eur. Phys. J. A \textbf{37}, 303 (2008). %[arXiv:0712.0354 [hep-ph]].

\bibitem{Sekihara:2015iha}
T.~Sekihara and E.~Oset,
%``Investigating the nature of light scalar mesons with semileptonic decays of D mesons,''
Phys. Rev. D \textbf{92}, 054038 (2015). %[arXiv:1507.02026 [hep-ph]].

% test: ===================
\bibitem{Stone:2013eaa}
S.~Stone and L.~Zhang,
%``Use of $B\to J/\psi f_0$ decays to discern the $q \bar{q}$ or tetraquark nature of scalar mesons,''
Phys. Rev. Lett. \textbf{111}, 062001 (2013). %[arXiv:1305.6554 [hep-ex]].

\bibitem{Cheng:2002ai}
H.~Y.~Cheng,
%``Hadronic D decays involving scalar mesons,''
Phys. Rev. D \textbf{67}, 034024 (2003). %[arXiv:hep-ph/0212117 [hep-ph]].

\bibitem{Wang:2009azc}
W.~Wang and C.~D.~Lu,
%``Distinguishing two kinds of scalar mesons from heavy meson decays,''
Phys. Rev. D \textbf{82}, 034016 (2010). %[arXiv:0910.0613 [hep-ph]].


% broken effects: ==========
\bibitem{Gronau:1995hm}
M.~Gronau, O.~F.~Hernandez, D.~London and J.~L.~Rosner,
%``Broken SU(3) symmetry in two-body B decays,''
Phys. Rev. D \textbf{52}, 6356 (1995). %[arXiv:hep-ph/9504326 [hep-ph]].

\bibitem{Wang:2023uea}
Y.~L.~Wang, H.~J.~Zhao and Y.~K.~Hsiao,
%``Topological SU(3)f approach for two-body {\ensuremath{\Omega}}c weak decays,''
Phys. Rev. D \textbf{111}, 016022 (2025). %[arXiv:2310.18896 [hep-ph]].

\end{thebibliography}

\documentclass[preprint,preprintnumbers, prd, floatfix, superscriptaddress,nofootinbib] {revtex4-1}

\usepackage{epsfig}
\usepackage{subfigure}
\usepackage{dcolumn}
\usepackage{bm}
\usepackage[usenames ,dvipsnames]{xcolor}
\usepackage{slashed}
\usepackage{graphicx,color}
\usepackage{hyperref}

\begin{document}
\title{Probing the elusive $\kappa/K_0^*(700)$ resonance in semileptonic $D$ decays}

\author{Yu-Kuo Hsiao}
\email{yukuohsiao@gmail.com}
\affiliation{School of Physics and Electronic Engineering, 
Shanxi Normal University, Taiyuan 030031, China}

\author{Yan-Li~Wang}
\email{ylwang0726@163.com}
\affiliation{School of Physics and Electronic Engineering, Shanxi Normal University,
Taiyuan 030031, China}

\author{Wen-Juan Wei}
\email{snza12321@163.com}
\affiliation{School of Physics and Electronic Engineering, 
Shanxi Normal University, Taiyuan 030031, China}

\date{\today}

\begin{abstract}
The $\kappa/K_0^*(700)$ meson remains the most elusive among the light scalar resonances, 
with its presence in weak decays obscured by limited precision 
in branching fraction measurements. As a result, 
the true nature of the $\kappa$ remains difficult to explore.
Through a partial-wave analysis of the semileptonic decay $D^+ \to K^-\pi^+ e^+ \nu_e$, 
we extract ${\cal B}(D^+ \to \bar\kappa^0 e^+ \nu_e, \bar\kappa^0 \to K^-\pi^+) 
=(2.2 \pm 0.1) \times 10^{-3}$. Previously considered negligible, 
this contribution is now shown to dominate the observed s-wave branching fraction.
This reveals that clear evidence for the $\kappa$ in weak decays 
has long existed, but was misidentified as part of the non-resonant background.
The extracted $D^+ \to \bar\kappa^0$ form factor, $f^+(0) = 0.32 \pm 0.01$, 
is significantly smaller than the $q\bar q$ prediction of $0.82 \pm 0.05$, 
and closely aligns with the $q^2\bar q^2$ expectation of $0.36 \pm 0.02$. Notably, 
this finding supports a compact tetraquark interpretation of the $\kappa$ meson.
\end{abstract}

\maketitle
\section{introduction}
Light scalar resonances, including the $f_0/f_0(980)$, $a_0/a_0(980)$, $\sigma/f_0(500)$, 
and $\kappa/K_0^*(700)$, have long posed significant challenges in identification.
For example, the decays $a_0 \to \pi\eta, K\bar K$ and $f_0 \to \pi\pi, K\bar K$ 
result in multiple competing channels, obscuring the resonance signals in scattering experiments.
The $\sigma$ and $\kappa$ states, in particular, exhibit broad decay widths 
comparable to their masses, resulting in diffuse or overlapping signals 
rather than distinct peaks in invariant mass spectra~\cite{Pelaez:2015qba}. 
Among them, the $\kappa$ remains the most elusive.  
Its identification as the final member of the light scalar nonet 
was only recently confirmed through high-precision dispersive analyses 
of $\pi K \to \pi K$ and $\pi\pi \to K\bar{K}$ scattering~\cite{pdg,Pelaez:2020uiw,Pelaez:2020gnd}.
Yet, despite numerous experimental efforts~\cite{E791:2002xlc,
FOCUS:2007mcb,CLEO:2008jus,
BESIII:2014oag,FOCUS:2005iqy,CLEO:2010enr,BaBar:2010vmf,
BESIII:2015hty,BESIII:2018jjm,BESIII:2024awg,BESIII:2024xjf,BESIII:2025fso},
no unambiguous evidence of the $\kappa$ has been established in weak decays. 
Limited data continue to hinder efforts to uncover the internal nature of the $\kappa$ meson.

The semileptonic decay $D \to K\pi \ell\nu_\ell$,
proceeding via the intermediate kaonic resonances $K_J \to K\pi$,
offers a clean environment to probe the $\kappa$ resonance in weak decays. 
Here, $\ell=e$ or $\mu$, and $K_J$ include $\kappa/K^*(700)$, $K^*/K_0^*(892)$, $K^*(1410)$, 
$K^*_0(1430)$, $K^*_2(1430)$~\cite{BaBar:2010vmf}.
Partial-wave analyses consistently show that the $K^*$ dominates,
accounting for approximately 95\% of the rate through the p-wave component~\cite{FOCUS:2005iqy,CLEO:2010enr,BaBar:2010vmf,BESIII:2015hty,
BESIII:2018jjm,BESIII:2024awg,BESIII:2024xjf}. 
Contributions from higher resonances are negligible, with
${\cal B}(D^+\to \bar K^*(1410)^0 e^+\nu_e,\bar K^*(1410)^0\to K^-\pi^+)
=(0\pm 0.5)\times 10^{-5}$ and 
${\cal B}(D^+\to \bar K_2^*(1430)^0 e^+\nu_e$, $\bar K_2^*(1430)^0\to K^-\pi^+)
=(3.7\pm 2.5)\times 10^{-5}$~\cite{BESIII:2015hty,pdg}, compared to 
the total branching ratio of $(3.77\pm 0.03\pm 0.08)\times 10^{-2}$~\cite{BESIII:2015hty}.
While the s-wave component accounts for about 5\% of the total rate, 
the scalar $K_0^*(1430)$ contributes less than 0.64\%~\cite{FOCUS:2005iqy}.
The remaining s-wave strength has, however, been attributed to a non-resonant scalar component, 
with the possible contribution from the $\kappa$ resonance systematically excluded~\cite{FOCUS:2005iqy,CLEO:2010enr,BaBar:2010vmf,BESIII:2015hty,
BESIII:2018jjm,BESIII:2024awg,BESIII:2024xjf,BESIII:2025fso}.

Given its scalar nature, any $\kappa$ contribution in $D \to K\pi \ell \nu_\ell$ 
must manifest in the s-wave channel and thus warrants careful investigation.  
To address this, we perform a complete and comprehensive partial-wave analysis 
of semileptonic $D\to K\pi e \nu_e$ decay, incorporating related non-leptonic decays such as
$D\to K\pi\pi,K\pi\rho$, to isolate and quantify the $\kappa$ contribution. 
Our results provide the first clear evidence of the $\kappa$ resonance 
in weak decays and offer critical test of its internal structure, 
including the possibility of a compact tetraquark configuration~\cite{Jaffe:1976ig,
Jaffe:1976ih,Close:2002zu,Pelaez:2003dy,Maiani:2004uc,Amsler:2004ps,
Jaffe:2004ph,tHooft:2008rus,Fariborz:2009cq,Weinberg:2013cfa,Agaev:2018fvz}.

\section{Formalism}
The four-body semileptonic $D\to K\pi e \nu_e$ decay 
proceeds via the quark-level transition $c\to s e \nu_e$. 
According to the effective Hamiltonian formalism~\cite{Buchalla:1995vs,Buras:1998raa},
the decay amplitude can be expressed as~\cite{BaBar:2010vmf,Hsiao:2023qtk} 
\begin{eqnarray}\label{amp1}
{\cal M} 
&=&\frac{G_F}{\sqrt 2}V_{cs}
\langle K\pi|(\bar s c)_{V-A}|D\rangle(\bar u_\nu v_e)_{V-A}\,,
\end{eqnarray}
where $(\bar a b)_{V-A}\equiv \bar a\gamma_\mu(1-\gamma_5)b$,
$G_F$ is the Fermi constant, and $V_{cs}$ 
the Cabibbo-Kobayashi-Maskawa (CKM) matrix element. 
The hadronic matrix elements include both resonant (RE) and non-resonant (NR) components~\cite{Lee:1992ih,Tsai:2021ota}
\begin{eqnarray}\label{FF1}
&&
\langle K \pi|(\bar s c)_{V-A}|D\rangle_{\rm RE}
=\langle K\pi|K_J\rangle D_{K_J}^{-1}\langle K_J|(\bar s c)_{V-A}|D\rangle\,,\nonumber\\
&&
\langle K \pi|(\bar s c)_{V-A}|D\rangle_{\rm NR}
=
h\epsilon_{\mu\alpha\beta\gamma}p_D^\alpha (p_K+p_{\pi})^\beta (p_K-p_{\pi})^\gamma
\nonumber\\
&&
+ir q_\mu+iw_+ (p_K+p_{\pi})_\mu+iw_-(p_K-p_{\pi})_\mu\,,
\end{eqnarray}
where $q_{\mu}=(p_D-p_K-p_{\pi})_\mu$.
The resonant term describes the process $D\to K_J$,
followed by resonance propagation $D_{K_J}^{-1}$, and decay into $K\pi$,
encoded in $\langle K\pi|K_J\rangle$. The form factors
$F_{\rm NR}=(h, r, w_{\pm})$ parameterize the non-resonant $D\to K\pi$ transition.
As noted in the introduction, the dominant resonance 
in semileptonic decay is the $K^*$, while the $\kappa$ 
is included specifically for our investigation. 
Contributions from higher resonances have been tested and found to be negligible. 
We therefore focus on the hadronic matrix elements for  
the $D$ to $K^*$ and $\kappa$ transitions,
given by~\cite{Cheng:2022vbw,Hsiao:2023qtk,Melikhov:2000yu} 
\begin{eqnarray}\label{FF2}
&&
\langle K^*|(\bar s c)_{V-A}|D\rangle=\epsilon_{\mu\alpha\beta\gamma}
\varepsilon^{\alpha}p_D^{\beta}p_{K^*}^{\gamma}\frac{2V}{m_D+m_{K^*}}
-i\bigg[\varepsilon_\mu
-\frac{\varepsilon\cdot p_D}{q^2}q_\mu\bigg](m_D+m_{K^*})A_1\,\nonumber\\
&&-i\frac{\varepsilon\cdot p_D}{q^2}q_\mu(2m_{K^*})A_0
+i\bigg[(p_D+p_{K^*})_\mu-
\frac{m^2_D-m^2_{K^*}}{q^2}q_\mu \bigg](\varepsilon\cdot p_D)\frac{A_2}{m_D+m_{K^*}}\;,
\nonumber\\
&&
\langle \kappa|(\bar s c)_{V-A}|D\rangle=
i{f^+(p_D+p_\kappa)_\mu+i(f^0-f^+)\frac{m^2_D-m^2_\kappa}{q^2}q_\mu}\,,
\end{eqnarray}
where $\varepsilon_\mu$ is the polarization vector of the $K^*$, and
$(V,A_0,A_1,A_2)$ and $f^{0,+}$ the associated form factors 
for the $K^*$ and $\kappa$, respectively.
The corresponding resonance propagators are given by
$D_{K^*}={m_{K^*}^2-t-i\,m_{K^*}\Gamma_{K^*}}(t)$ and $D_{\kappa}
={m_{\kappa}^2-t-\Gamma_{\kappa}(t)^2/4-i\,m_{\kappa}
\Gamma_{\kappa}}(t)$~\cite{BaBar:2010vmf,Cheng:2022vbw}.
The form of $D_{\kappa}$ is appropriate for a broad resonance with $\Gamma \sim m$~\cite{pdg}.
The energy-dependent widths are defined as
$\Gamma_{K^*}(t)=\Gamma_{K^*}^0[q^3 m_{K^*}X_1^2(q)]/[q_0^3 \sqrt t X_1^2(q_0)]$ and 
$\Gamma_{\kappa}(t)
=\Gamma_{\kappa}^0(q m_{\kappa})/(q_0 \sqrt t)$~\cite{BaBar:2010vmf,Cheng:2022vbw},
where $\Gamma_{K_J}^0=\Gamma_{K_J}(m_{K_J}^2)$ is the on-shell width,
and $q=|\vec{p}_K|=|\vec{p}_\pi|$ is the three-momentum of the decay products
in the $K_J$ rest frame. The quantity $q_0$ corresponds to $q$ evaluated at $t=m_{K_J}^2$.
The Blatt-Weisskopf barrier factor is defined as $X_1^2(z)\equiv [(z r_{BW})^2+1]^{-1}$, 
with $r_{BW}\simeq 3.0$~GeV$^{-1}$~\cite{BESIII:2015hty}.
The hadronic matrix elements for the strong decays $K^*\to K\pi$ and $\kappa \to K\pi$ 
are given by $\langle K\pi|K^*\rangle=g_{K^*K\pi} \varepsilon\cdot (p_K-p_\pi)$ and 
$\langle K\pi|\kappa\rangle=g_{\kappa K\pi}$~\cite{Cheng:2022vbw,Cheng:2020ipp}, 
where $g_{K^* K\pi}$ and $g_{\kappa K\pi}$ are the corresponding strong coupling constants.

To perform the partial-wave analysis, 
we adopt the expression for the differential decay width~\cite{Lee:1992ih}
\begin{eqnarray}\label{Gamma1}
d\Gamma
&=&\frac{1}{3}\frac{G_F^2|V_{cs}|^2}{(4\pi)^5 m_D^3}X\alpha_{\bf M}\alpha_{\bf L}
\bigg\{2|F_{10}|^2+\frac{2}{3}\bigg[
|F_{11}|^2+|F_{21}|^2+|F_{31}|^2\bigg]\bigg\}dsdt\,,
\end{eqnarray}
where the kinematic factors are defined as 
$X=[(m_D^2-s-t)^2/4-st]^{1/2}$,
$\alpha_{\bf M}=\lambda^{1/2}(t,m_K^2,m_\pi^2)/t$, and
$\alpha_{\bf L}=\lambda^{1/2}(s,m_\ell^2,m_{\nu_\ell}^2)/s$,
with $s=(p_\ell+p_{\nu_\ell})^2$, and the K\"all\'en function defined as
$\lambda(a,b,c)=a^2+b^2+c^2-2ab-2bc-2ca$. 
The partial-wave amplitudes $F_{ij}$ include the s-wave component $F_{10}$ 
and p-wave components $F_{11}$, $F_{21}$, and $F_{31}$, 
given by
\begin{eqnarray}\label{Fij}
&&
F_{10}=X \bigg[\bigg(w_+ +\frac{m_K^2-m_\pi^2}{t}w_-\bigg)+2\hat f^+\bigg]\,,\nonumber\\
&&
F_{11}=\alpha_{\bf M}\bigg\{\frac{1}{2}(m_D^2-s-t)[w_- +(m_D+m_{K^*})\hat A_1]
-2X^2\frac{\hat A_2}{m_D+m_{K^*}}\bigg\}\,,\nonumber\\
&&
F_{21}=\alpha_{\bf M}\sqrt{2 st}[w_- +(m_D+m_{K^*})\hat A_1]\,,\;
\nonumber\\
&&
F_{31}=\alpha_{\bf M}X\sqrt{2 st}\bigg(h-\frac{2\hat V}{m_D+m_{K^*}}\bigg)\,,
\end{eqnarray}
with the intermediate form factors defined as 
$(\hat V,\hat A_1,\hat A_2)=g_{K^* K\pi}D^{-1}_{K^*}\times (V,A_1,A_2)$
and $\hat f^+=g_{\kappa K\pi}D^{-1}_{\kappa}\times f^+$.
Notably, this framework incorporates both non-resonant and $\kappa$-resonant contributions 
for the first time, enabling a more complete and systematic treatment of 
both s- and p-wave components in semileptonic decays.

Currently, the partial analyses are insufficient to determine additional form factors.
In this context, non-leptonic decays such as $D\to K\pi\pi$ and $D\to K\pi\rho$,
which involve in the same $D\to K\pi$ transition, can serve as complementary probes.
Notably, experimental analyses can exclude resonant $K^*$ contributions, 
allowing for a cleaner investigation of non-resonant and $\kappa$-resonant form factors.
The corresponding decay amplitudes can be expressed as~\cite{Escribano:2023zjx,Boito:2009qd}
\begin{eqnarray}\label{amp2}
{\cal M}_{\pi,\rho}\simeq\frac{G_F}{\sqrt 2}V_{cs}^*V_{ud}
a_1\langle \pi,\rho|(\bar u d)_{V-A}|0\rangle
\langle K\pi|(\bar s c)_{V-A}|D\rangle\,,
\end{eqnarray}
where the color-suppressed internal $W$-boson emission contributions are neglected.
The vacuum-to-meson matrix elements are given by
$\langle\pi,\rho|(\bar u d)_{V-A}|0\rangle=(if_\pi p_\pi^\mu,m_\rho f_\rho \varepsilon^\mu)$,
where $f_{\pi(\rho)}$ denotes the decay constant.
The coefficient $a_1=c_1+c_2/N_c$ arises from the factorization approach~\cite{Ali:1998eb}, 
where $c_1$ and $c_2$ are scale-dependent Wilson coefficients,
and $N_c$ is the color number.
Using the standard equation: $d\Gamma_{\pi,\rho}=1/[(2\pi)^3 32 M_D^3]
|{\cal M}_{\pi,\rho}|^2 dm_{12}^2 dm_{23}^2$~\cite{pdg},
one can compute the decay widths for $D\to K\pi\pi$ and $D\to K\pi\rho$, 
where $m_{12}^2=(p_1+p_2)^2$ and $m^2_{23}=(p_2+p_3)^2$ are
the invariant mass combinations. The final-state momenta 
$p_1$, $p_2$, and $p_3$ correspond to $K$, $\pi$, and $\pi(\rho)$, respectively.

\section{Numerical analysis}
In our numerical analysis, we adopt the Wolfenstein parameterization 
for the CKM matrix elements, taking $V_{cs}=V_{ud}=1-\lambda^2/2$
with $\lambda=0.22453\pm 0.00044$~\cite{pdg}. 
The decay constants used in Eq.~(\ref{amp2}) are 
$(f_\pi,f_\rho)=(130.4\pm 0.2,210.6\pm 0.4)$~MeV~\cite{pdg,Hsiao:2014mua}.
For the Wilson coefficients,
we use $(c_1,c_2)=(1.26,-0.51)$ at the $m_c$ scale~\cite{Buchalla:1995vs}.
Within the generalized factorization framework~\cite{Ali:1998eb}, 
and by varying the (effective) color number $N_c^{(eff)}=(2,3,\infty)$,
we obtain $a_1=1.1\pm 0.1$. The quoted 10\% uncertainty 
reflects  the modest sensitivity of $a_1$ to 
non-factorizable QCD loop corrections~\cite{Wang:2024mjw}.

Partial-wave analyses of the decay $D\to K\pi e \nu_e$ 
conventionally parameterize the form factors $V$, $A_1$, and $A_2$ 
using single-pole forms~\cite{BaBar:2010vmf, BESIII:2015hty, 
BESIII:2018jjm, BESIII:2024awg, BESIII:2024xjf,BESIII:2025fso}:
\begin{eqnarray}\label{VandA}
V(q^2)=\frac{V(0)}{1-\frac{q^2}{m_V^2}}\,,\;
A_1(q^2)=\frac{A_1(0)}{1-\frac{q^2}{m_A^2}}\,,\;
A_2(q^2)=\frac{A_2(0)}{1-\frac{q^2}{m_A^2}}\,,
\end{eqnarray}
where $m_V$ and $m_A$ denote the vector and axial-vector pole masses, respectively.
The scalar form factor $f^+(q^2)$, 
along with the non-resonant vector-type form factors $w_\pm(q^2)$,
and the tensor-type form factor $h(q^2)$, 
are modeled following Ref.~\cite{Hsiao:2023qtk} as
\begin{eqnarray}\label{fwh}
f^+(q^2)=\frac{f(0)}{(1-\frac{q^2}{m_D^2})^n}\,,\;
w_\pm(q^2)=\frac{w_\pm(0)}{(1-\frac{q^2}{m_D^2})^n}\,,\;
h(q^2)=\frac{h(0)}{(1-\frac{q^2}{m_D^2})^n}\,,\;
\end{eqnarray}
where $n = 2$ reflects a double-pole behavior.
The form factors $A_0$, $r$ and $f^0$ 
are omitted from Eqs.~(\ref{VandA}) and (\ref{fwh}), 
as they contribute negligibly to semi- and non-leptonic amplitudes.
Using isospin symmetry, the non-resonant form factors
$F_{\rm NR}$ for the $D^+\to K^-\pi^+$ transition are 
four (two) times as large as those for $D^+\to K^0_S\pi^0 (D^0\to K^-\pi^0)$.
Using the strong decay amplitudes and PDG values for meson masses and widths~\cite{pdg}, 
we extract the strong coupling constants as $g_{\bar K^{*0} K^-\pi^+}=2g_{\bar K^{*0} K^0_S\pi^0}=4.5$ and
$g_{\bar\kappa^0 K^-\pi^+}=2g_{\bar \kappa^0 K^0_S\pi^0}=4.7$~GeV.

We perform a minimum $\chi^2$-fit using the equation
$\chi^2=\Sigma_i[({\cal B}^i_{\rm th}-{\cal B}^i_{\rm ex})/\delta{{\cal B}^i_{\rm ex}}]^2$
$+\Sigma_j[(F_{\rm th}^j-F_{\rm ex}^j)(\delta F_{\rm ex}^j)]^2$,
where ${\cal B}_{\rm th}^i$ are the theoretical branching fractions calculated from
the decay amplitudes for $D\to K\pi\ell\nu_\ell$ from Eq.~(\ref{amp1}),
and $D\to K\pi\pi$ and $D\to K\pi\rho$ from Eq.~(\ref{amp2}). 
The corresponding experimental inputs ${\cal B}_{\rm ex}^i$ and their uncertainties
are summarized in Table~\ref{data1}. The second term incorparate the form factors
$F^j=V(0)$, $A_1(0)$, $A_2(0)$, with theoretical and experimental values
denoted by $F^j_{\rm th}$ and $F^j_{\rm ex}$, respectively.
We adopt the experimental inputs from Ref.~\cite{BESIII:2015hty}:
$(V(0),A_1(0),A_2(0))=(0.83\pm 0.05,0.59\pm 0.02,0.47\pm 0.03)$,
with the pole masses $(m_V,m_A)=(2.0,2.6)$~GeV.
We assume $w(0)=w_\pm(0)$ following model calculations of meson transitions, 
which suggests that the timelike and spacelike form factors are nearly equal~\cite{Melikhov:2000yu}.
Treating $f(0)$, $w(0)$, and $h(0)$ in Eq.~(\ref{fwh}) as free parameters,
the $\chi^2$-fit yields
\begin{eqnarray}\label{fit_result}
f(0)=0.32\pm 0.01\,,\;
w(0)=(0.94\pm 0.14)~{\rm GeV}^{-1}\,,\;
h(0)=(4.24\pm 0.79)~{\rm GeV}^{-3}\,,
\end{eqnarray}
with $\chi^2/\text{n.d.f.} =1.2$, 
where $\text{n.d.f.} = 2$ denotes the number of degrees of freedom.  
Using the fitted values in Eq.~(\ref{fit_result}), 
we compute the branching fractions listed as our work in Table~\ref{data1}.

%
\begin{table}[b]
\caption{Branching fractions of $D\to K\pi e\nu_e$, $D\to K\pi\pi$,
and $D\to K\pi\rho$. The values with the star notations are
the experimental inputs in the global fit. In the second column of our work,
the first uncertainties incorporate those of the form factors in Eq.~(\ref{fit_result}),
and the second one, specifically in non-leptonic decays, is from $\delta a_1$.}\label{data1}
{
\tiny
\begin{tabular}{|ccc|}
\hline
branching fraction&our work&data\\
\hline\hline
$10^{2}{\cal B}(D^+\to K^-\pi^+ e^+\nu_e)$
&$3.7\pm0.1$
&$(3.77\pm 0.09)^*$~\cite{BESIII:2015hty}\\
$10^{3}{\cal B}(D^+\to (K^-\pi^+)_{\rm s-wave} e^+\nu_e)$
&$2.3\pm 0.1$
&$(2.28\pm 0.11)^*$~\cite{pdg,BESIII:2015hty}\\
$10^2 {\cal B}(D^+\to \bar K^{*0} e^+\nu_e,\bar K^{*0}\to K^-\pi^+)$
&$3.5\pm 0.1$
& $3.54\pm 0.09$~\cite{BESIII:2015hty}\\
$10^3 {\cal B}(D^+\to \bar\kappa^0 e^+\nu_e,\bar\kappa^0\to K^-\pi^+)$
&$2.2\pm 0.1$
& ignored~\cite{BESIII:2015hty}\\
$10^3 {\cal B}(D^+\to (K^-\pi^+)_{\rm NR} e^+\nu_e)$
&$0.14\pm 0.04$
&$2.28\pm 0.11$~\cite{BESIII:2015hty}\\
\hline\hline
$10^2{\cal B}(D^+\to\pi^+ (K^0_S\pi^0)_{\rm NR}+\pi^+(\bar \kappa^0\to) K^0_S\pi^0)$
&$1.2\pm 0.3\pm 0.2$
&$(1.37\pm 0.40)^*$~\cite{BESIII:2014oag,pdg}\\
$10^3{\cal B}(D^+\to \pi^+ (K^0_S\pi^0)_{\rm NR})$
&$9.1\pm 3.0\pm 1.8$
&$3\pm 4$~\cite{BESIII:2014oag,pdg}\\
$10^3{\cal B}(D^+\to \pi^+ \bar \kappa^0,\bar \kappa^0\to K^0_S\pi^0)$
&$2.4\pm 0.1\pm 0.5$
&$6^{+5}_{-4}$~\cite{BESIII:2014oag,pdg}\\
$10^{2}{\cal B}(D^0\to \pi^+ (K^-\pi^0)_{\rm NR})$
&$1.4\pm 0.5\pm 0.3$
&$(1.15\pm 0.60)^*$~\text{\cite{pdg}}\\
$10^{3}{\cal B}(D^+\to \rho^+ (K^0_S\pi^0)_{\rm NR})$
&$4.8\pm 1.3\pm 1.0$
&$(4.8\pm 0.5)^*$~\text{\cite{pdg}}\\
$10^2{\cal B}(D^+\to \pi^+ (K^-\pi^+)_{\rm NR})$
&$3.6\pm 1.2\pm 0.7$
&---\\
$10^2{\cal B}(D^+\to \pi^+ \bar \kappa^0,\bar \kappa^0\to K^-\pi^+)$
&$1.0\pm 0.1\pm 0.2$
&---\\
\hline
\end{tabular}}
\end{table}
%

\section{Discussions and Conclusion}
As shown in Table~\ref{data1},
the five data points used in the global fit are well reproduced with the fitted form factors,
reflecting the goodness of the fit  with $\chi^2/n.d.f.\simeq 1$. 
Ref.~\cite{BESIII:2014oag} discovers that $D\to \pi(K\pi)_{\rm s-wave}$ channel 
receives contributions from three sources: 
a non-resonant component, the $K_0^*(1430)$, and the $\kappa$. 
Based on this decomposition, we predict ${\cal B}^{00}_{{\rm NR},\kappa}
\equiv{\cal B}(D^+\to \pi^+ (K^0_S\pi^0)_{\rm NR},\pi^+(\bar \kappa^0\to) K^0_S\pi^0)$ and
${\cal B}^{-+}_{{\rm NR},\kappa}$, shown in Table~\ref{data1}.
The latter is obtained by replacing $K^0_S\pi^0$ with $K^-\pi^+$ in ${\cal B}^{00}_{{\rm NR},\kappa}$, 
taking into account the identical final-state pions in the calculation.
Including the measured contributions from $\bar K_0^*(1430)^0$ in the PDG~\cite{pdg}, 
${\cal B}^{00}_{K_0}=(2.7\pm 0.9)\times 10^{-3}$ and 
${\cal B}^{-+}_{K_0}=(1.25\pm 0.06)\times 10^{-2}$, we obtain total s-wave branching fractions
${\cal B}_{\rm NR}^{00}+{\cal B}_{\kappa}^{00}+{\cal B}_{K_0}^{00}=(1.4\pm 0.4)\times 10^{-2}$ 
and ${\cal B}_{\rm NR}^{-+}+{\cal B}_{\kappa}^{-+}+{\cal B}_{K_0}^{-+}=(6.0\pm 1.4)\times 10^{-2}$.
These results are consistent with the experimental values
${\cal B}(D^+\to\pi^+(K^0_S\pi^0)_{\rm s-wave})
=(1.27\pm 0.33)\times 10^{-2}$~\cite{BESIII:2014oag,pdg} and 
${\cal B}(D^+\to\pi^+(K^-\pi^+)_{\rm s-wave})
=(7.52\pm 0.17)\times 10^{-2}$~\cite{pdg},
further supporting the validity of the extracted form factors.

In semileptonic $D$ decays, we predict
${\cal B}^{\rm semi}_{\kappa}\equiv
{\cal B}(D^+\to \bar\kappa^0 e^+\nu_e,\bar\kappa^0\to K^-\pi^+)$
and ${\cal B}^{\rm semi}_{\rm NR}\equiv {\cal B}(D^+\to (K^-\pi^+)_{\rm NR} e^+\nu_e)$, 
with the non-resonant component further decomposed into s- and p-wave contributions
${\cal B}^{\rm semi}_{\rm NR(S)}=(5.9\pm 2.0)\times 10^{-5}$ and
${\cal B}^{\rm semi}_{\rm NR(P)}=(5.5\pm 1.8)\times 10^{-5}$.
Our results show that ${\cal B}^{\rm semi}_{\kappa}$ exceeds
${\cal B}^{\rm semi}_{\rm NR(S)}$ by an order of magnitude, establishing the $\kappa$
as the dominant source of the s-wave branching fraction. Crucially, our analysis also reveals,
for the first time, a non-zero p-wave contribution from the non-resonant amplitude.
These findings challenge the prevailing experimental assumption~\cite{FOCUS:2005iqy,BaBar:2010vmf,BESIII:2015hty,
BESIII:2018jjm,BESIII:2024awg,BESIII:2024xjf,BESIII:2025fso}, 
which holds that contributions from both the s-wave $\kappa$
and p-wave non-resonant terms are negligible, 
attributing the observed s-wave strength
primarily to a non-resonant background.

The form factors $h$ and $w_\pm$ successfully accounts for  
${\cal B}(D \to K\pi\pi,K\pi\rho)$. Our baseline assumes $w_+ = w_-$, 
motivated by model-based estimates suggesting that the timelike and spacelike
form factors are nearly equal, warranting examination.
Indeed, this assumption has been tested against three alternatives: 
(i) treating $w_+$ and $w_-$ as independent parameters, 
(ii) imposing $w_+ = -w_-$, and (iii) setting either $w_+ = 0$ or $w_- = 0$. 
None yielded acceptable fits, producing large $\chi^2$ or unstable extractions of $h(0)$. 
Nonetheless, all scenarios consistently predict
${\cal B}^{\rm semi}_{\rm NR(S)}\sim$ a few $\times 10^{-5}$
and ${\cal B}^{\rm semi}_{\rm NR(S)}\simeq {\cal B}^{\rm semi}_{\rm NR(P)}$,
confirming that the non-resonant s-wave component alone cannot account for 
the total observed s-wave strength. 

The $\kappa$ meson remains the most elusive of the light scalar states, 
with its existence in weak decays long considered inconclusive. 
However, the precise extraction of $f^+(0)$,
remarkably stable across various theoretical scenarios,
allows for reliable predictions of the branching fractions 
${\cal B}^{00,-+,{\rm semi}}_{\kappa}$, treating the $\kappa$ as a genuine resonance. 
In particular, our result for ${\cal B}^{\rm semi}_{\kappa}$, which dominates the s-wave contribution, 
reveals that clear evidence for the $\kappa$ in weak decays 
has existed for some time but was misidentified as non-resonant background.

Our investigation provides new insight into the internal nature of the light scalar mesons.
In semileptonic decays $D\to M_1 M_2 \ell\nu_\ell$, 
only the s-wave $M_1 M_2$ pair can undergo s-wave scattering, 
potentially forming a molecular state~\cite{Branz:2007xp,Sekihara:2015iha,Hsiao:2023qtk}. 
However, the small branching fraction ${\cal B}^{\rm semi}_{\rm NR(S)} \sim 10^{-5}$ 
indicates that such meson-meson scattering, and hence a $\kappa$-molecule formation, 
is highly suppressed. In contrast, treating the $\kappa$ as a compact state 
leads to a successful description of the data via the form factor $f^+(q^2)$.

Regarding $\bar \kappa^0$ as a compact state, 
it may correspond to either a conventional $q\bar q$ meson~($|s\bar d\rangle$) 
or a $q^2\bar q^2$ tetraquark ($|s\bar d u\bar u\rangle$)~\cite{Maiani:2004uc,Stone:2013eaa}.
Likewise, the light scalar mesons $S_n=|(u\bar u + d\bar d)/\sqrt{2}\rangle$ and 
$S_{ns}=|(u\bar u + d\bar d)s\bar s/\sqrt{2}\rangle$ 
represent $q\bar q$ and $q^2\bar q^2$ configurations, respectively, 
and mix to form physical states such as $f_0$, $\sigma_0$, and $a_0^0$. 
Under $SU(3)$ flavor $[SU(3)_f]$ symmetry~\cite{Stone:2013eaa,
Cheng:2002ai,Wang:2009azc,Hsiao:2023qtk},
the corresponding $D^+ \to \bar\kappa^0$ form factors are related by
$F_{q\bar q}^{D^+\to\bar\kappa^0}=\sqrt{2}(f_K/f_\pi) F^{D^+\to S_n}$ and 
$F_{q^2\bar q^2} = \sqrt{2}(f_\pi/f_K) F^{D^+\to S_{ns}}$, 
with $f_K/f_\pi = 1.2$ and $f_\pi/f_K = 0.8$ accounting for $SU(3)_f$ breaking~\cite{Gronau:1995hm,Wang:2023uea}.
Using $F^{D^+\to S_n} = 0.47\pm 0.02$ and 
$F^{D^+\to S_{ns}} = 0.31\pm 0.02$~\cite{Hsiao:2023qtk},
we obtain $F_{q\bar q}^{D^+\to\bar\kappa^0}=0.82 \pm 0.05$ and 
$F_{q^2\bar q^2}^{D^+\to\bar\kappa^0}=0.36 \pm 0.02$. 
The extracted value $f^+(0) = 0.32 \pm 0.01$ lies far away from the $q\bar q$ expectation  
and much closer to the tetraquark prediction, providing strong evidence 
for a $q^2\bar q^2$ configuration of the $\kappa$ resonance.

In summary, we have performed a partial-wave analysis of the semileptonic decay 
$D \to K\pi e \nu_e$, complemented by an investigation of the nonleptonic decays 
$D \to K\pi\pi$ and $D \to K\pi\rho$. Using non-resonant $D \to K\pi$ form factors 
$h(q^2)$ and $w_\pm(q^2)$, along with the $D \to \kappa$ form factor $f^+(q^2)$,
all extracted via a global fit, we determined the branching fractions:
${\cal B}^{\rm semi}_{\rm NR(S)}\equiv
{\cal B}_{\rm NR}(D^+\to (K^-\pi^+)_{\rm s-wave} e^+\nu_e)=(5.9\pm 2.0)\times 10^{-5}$,
${\cal B}^{\rm semi}_{\rm NR(P)}\equiv
{\cal B}_{\rm NR}(D^+\to (K^-\pi^+)_{\rm p-wave} e^+\nu_e)=(5.5\pm 1.8)\times 10^{-5}$,
and ${\cal B}^{\rm semi}_{\kappa}\equiv
{\cal B}(D^+\to \bar\kappa^0 e^+\nu_e,\bar\kappa^0\to K^-\pi^+)=(2.2\pm 0.1)\times 10^{-3}$.
These results clearly indicate that the $\kappa$ resonance dominates the s-wave branching fraction, 
overturning the long-standing interpretation in the literature, which assumed that 
both ${\cal B}^{\rm semi}_{\kappa}$ and ${\cal B}^{\rm semi}_{\rm NR(P)}$ were negligible, 
and that the s-wave strength originated mainly from the non-resonant component
${\cal B}^{\rm semi}_{\rm NR(S)}$.
Our precise determination of ${\cal B}^{\rm semi}_\kappa$ indicates that 
first evidence for the $\kappa$ in weak decays 
has long been present but misidentified as non-resonant background.
The extracted form factor $f^+(0) = 0.32 \pm 0.01$ 
significantly deviates from the value expected for a conventional $q\bar{q}$ state ($0.82 \pm 0.05$), 
and instead agrees with predictions for a compact tetraquark configuration ($0.36 \pm 0.02$). 
This notable agreement provides strong support for interpreting light scalar mesons as tetraquark states.

\section*{ACKNOWLEDGMENTS}
This work was supported in parts by NSFC (Grants No.~12175128 and No.~11675030).

\end{document}